# Advancing Multi-Modal Sensing through Expandable Modality Alignment

### v1

@@ 3. Babel Overview
The crux of this design lies in the efficient alignment of representations across pre-trained encoders. Thereby, we introduce the modality tower, consisting of the pre-trained encoder, and the concept alignment module. The encoder could be based on signal processing and neural networks from existing DL models. The concept alignment module then aligns embeddings (features) from encoders. During training, pre-trained encoders are frozen, and the concept alignment module is updated.
In the design of the expandable model architecture, we try to convert the contrastive training process with N-tuple samples into a sequence of training phases involving only paired samples, thereby significantly reducing the need for tupled samples, rendering the alignment of multiple modalities truly feasible.
@@ 4.1. Assembling Modality Towers
Comparing with conventional modality alignment methods i.e., CLIP, Babel’s key design lies in the utilization of a pre-trained encoder within a singular modality, proving particularly effective for sensing modalities.
@@
The particular encoder for each modality is chosen based on the following criteria. For modalities dedicated to sensing tasks, such as mmWave, we tend to choose signal processing-based encoders, due to their capability to extract universally applicable features with well-defined physical meanings. For more ubiquitous sensing modalities, like Wi-Fi, which are often noisy, we lean towards DL-based encoders, owing to their proficiency in de-noising. As these DL encoders are trained on specific datasets, it is essential to avoid models with the domain shift during the selection. Therefore, our primary consideration is the model’s capacity. Eventually we evaluate and compare the selected candidates by fine-tuning and testing them on a variety of singular modality datasets. The encoder demonstrating superior generality is chosen.
To further boost the performance on certain modalities, we also propose our modality tower augmentation technique, which is elaborated upon in §4.3.
Figure 4. The alignment of two modality towers with pre-trained encoders and concept alignment modules.
@@ 4.2. Aligning Modality Towers
where $L_{\alpha\leftarrow\beta}^{M}$ and $L_{\beta\leftarrow\alpha}^{M}$ denote the computed contrastive loss transitioning from modality $\alpha$ to modality $\beta$ and vice versa within the batch $M$, as defined subsequently,
(4) $L_{\alpha\leftarrow\beta}^{M}=-\sum_{i=1}^{m}log(\frac{\exp(sim(P_{\alpha}^{i})/\tau)}{\sum\exp(sim(N_{\alpha}^{i})/\tau}),$
@@ 5.2. Growth Orders
Babel transforms the $N$-tuple modality alignment into a sequence of two-modality alignment phases, thereby raising a potential question regarding the differences between the conventional completed alignment and our expandable alignment with varying modality growth orders. To analyze this, we utilize a three-modality alignment, i.e., IMU, skeleton, and video from UTD-MHAD dataset (Chen et al., 2015), as an example.
As depicted in Fig.6, we utilize t-SNE to render the representation space of each modality visible. As evident in Figure 6(a), before alignment, features that have not undergone alignment training exhibit significant distribution differences. Fig. 6(b) shows the conventional triplet alignment successfully bridge the modality gaps, aligning the three modalities. In contrast, the expandable network architecture within Babel employs a sequence of two-modality alignment training phases as a replacement for the joint alignment. As illustrated in Fig. 6(c), we initially align the IMU and skeleton modalities followed by the video modality, effectively bridging the modality gaps as well.
@@ 6. Adaptive Training Strategy
We in further propose our training strategies to optimally integrate the insights derived from the newly aligned modality during network growth. Specifically, we implement two strategies for the training of the concept alignment module and the prototype network, respectively.
@@
Gradients effectively indicate how the loss function varies with the model parameters. If training pairs don’t provide new insights to the trunk network during the growth, it leads to small gradients in the junction modality tower, prompting a higher weight, bringing the branch network nearer to the trunk. When the branch modality tower’s gradients are significant, the assigned weight speeds up the absorption of insights from the trunk network, ensuring alignment in the unified representation space.
@@ 7.1. Data Preparation
Overall,we utilize five datasets for the alignment, as itemized in Table 1, comprising paired samples across divergent dual modalities. These datasets are for human activity recognition (HAR) tasks, but the certain activities are totally different. Despite the provision of activity labels within these datasets, we adopt a self-supervised training approach, labels are not used. Throughout every dataset, depth signals undergo a conversion into a human skeleton format. As such, we employ the term skeleton to denote the depth modality.
@@
Dataset Modalities # Pairs
UTD-MHAD (Chen et al., 2015) IMU and Skeleton $613$
MM-Fi (Yang et al., 2023c) LiDAR and Video $17,528$
OPERANet (Bocus et al., 2021) Wi-Fi and Skeleton $25,433$
XRF55 (Wang et al., 2024) mmWave and Wi-Fi $42,900$
Kinetics-400 (Kay et al., 2017) Video and Skeleton $234,619$
Skeleton and IMU pairs. UTD-MHAD dataset (Chen et al., 2015) is used, encompassing the skeleton and 9-axis IMU data pairs, captured via the Microsoft Kinect sensor and the wearable inertial sensor with respective sampling rates of 30Hz and 50Hz. The dataset embodies 27 distinct actions performed by 8 subjects. Each subject repeated the action for 4 times, totaling 861 paired samples. We use 613 pairs for the training.
@@
mmWave and Wi-Fi pairs. XRF55 dataset (Wang et al., 2024) is used, which is collected from a TI IWR6843ISK radar for mmWave and Intel 5300 for Wi-Fi CSI. It includes HAR data from 39 subjects performing 55 unique actions, each repeated 20 times. In total, 42,900 pairs are provided for training.
Video and skeleton pairs. Kinetics-400 dataset (Kay et al., 2017) is used, which contains 400 distinct human action classes, each characterized by at least 400 video clips extracted from YouTube. Each clip, approximately 10 seconds long, portrays a variety of human actions. The skeleton is extracted from use clips using OpenPose (Cao et al., 2017). Overall, as a dataset in vision modality, the Kinetics dataset provides 23,4619 pairs.
@@ 7.2. Data Augmentation
We implement two data augmentation techniques on the raw data, ultimately enlarging the data pairs by 600$\times$. (i) Down-sampling. Raw pairs undergo down-sampling at different ratios, simulating diverse sampling rates on various devices or accelerating the action at distinct ratios. This method augments the raw pairs by a factor of 300$\times$. (ii) Action-segmentation. The raw action sequence is randomly truncated, simulating incomplete activity sensing. We ensure the segmented sequence’s shortest length is over 50% of the original length. This method amplifies the raw pairs by a factor of 300$\times$.
7.3. Selections of Pre-trained Encoders
Next we introduce the pre-trained encoders we use for building the modality alignment network.
@@ 7.4. Training Details
We employ the AdamW optimizer (Loshchilov and Hutter, 2019) with a batch size of 256 and an initial learning rate of $1\times 10^{-4}$. For each phase of network growth, we judiciously allocate a varying number of training epochs, typically up to 500, or cease the training process once convergence is attained. The learning rate for downstream tasks is adjusted between 0.001 and 0.1. We train on two NVIDIA A100 GPUs, spending around 20 hours to align six modalities.
@@ 8. Evaluation
We evaluate pre-trained Babel by employing a typical downstream sensing task, human activity recognition (HAR). Furthermore, we would demonstrate two applications enabled by Babel, namely the cross-modality retrieval and the LLM integration.
@@ 8.1. Evaluation on HAR
We evaluate Babel on 8 datasets, including 4 multi-modal datasets namely UTD-MHAD (Chen et al., 2015), OPERANet (Bocus et al., 2021), XRF55 (Wang et al., 2024) and MM-Fi (Yang et al., 2023c), 4 singular-modal datasets, namely UCI (Reyes-Ortiz et al., 2016), Widar3.0 (Zhang et al., 2022), mRI (An et al., 2022a) and MSRAction3D (Li et al., 2010).
@@
We compare Babel with a broad range of baselines, including the multi-modal sensing baseline Cosmo (Ouyang et al., 2022), SOTA singular-modal sensing baselines, LIMU-BERT (Xu et al., 2021), SenseFi (Yang et al., 2023b), MARS (An and Ogras, 2021), MeteorNet (Liu et al., 2019), and PointTransformer (Zhao et al., 2020). We compare with MLLMs which hold potential for interpreting sensing signals, including OneLLM (Han et al., 2023) and M4 (Yuan et al., 2024).
Unless otherwise noted, the results for Babel are obtained from a one-shot setting, where only one labeled sample per class is used to train the downstream classifier. Given the inherent difficulty of obtaining labeled samples for sensing applications, this setting highlights Babel’s performance as a pre-trained network.
Table 2. One-shot HAR classification accuracy of each modality before and after the modality alignment using Babel on four multi-modal datasets.
Modality HAR Classes Dataset Before Align. After Align. Fusion IMU 27 UTD-MHAD (Chen et al., 2015) $20.19$% 31.77% 33.17% Video 27 UTD-MHAD (Chen et al., 2015) $19.23$% 21.35% Wi-Fi 6 OPERANet (Bocus et al., 2021) $23.15$% 33.89% 58.97% mmWave 55 XRF55 (Wang et al., 2024) $30.32$% 50.30% LiDAR 27 MM-Fi (Yang et al., 2023c) $28.43$% 43.91% - Skeleton 27 UTD-MHAD (Chen et al., 2015) $49.04$% 61.06%
8.1.1. Performance on Multi-modal Datasets.
@@
Table 2 shows the evaluation results on four multi-modal datasets. The modality alignment technique in Babel significantly improves performance in each individual modality, particularly in sensing modalities. For instance, classification accuracy for 27 human activities in the IMU modality increases from 20.19% to 31.77% before and after alignments. For the skeleton modality, there is a significant improvement by 12.02%. The Wi-Fi modality sees an approximate 10.74% enhancement. The mmWave modality shows a substantial increase from 30.32% to 50.30%, and the LiDAR modality achieves an accuracy of 43.91%, up from 28.43%. Overall, Babel brings around averaged 12% accuracy improvement on six aligned modalities across various datasets. Such gains are achieved by aligning each modality into a unified representation space, facilitating mutual learning. Sensing modalities benefit significantly, while gains are limited for video modalities.
The unified representation space in Babel allows for effective fusion, as shown in Table 2. When IMU and Video modalities are fused, Babel achieves a 33.17% accuracy on UTD-MHAD (Chen et al., 2015), outperforming both the individual IMU and video modalities. Likewise, a 58.97% accuracy is achieved on XRF55 (Wang et al., 2024) when Wi-Fi and mmWave modalities are merged. Note that even modality combinations (like IMU&Video fusion) not included in pre-training datasets are evaluated and obtain the superior performance, highlighting Babel’s flexibility and offering developers numerous opportunities to choose any one or combined modalities for their tasks.
8.1.2. Performance on Singular-modal Datasets.
@@
The supervised learning performance of Babel on four full singular-modal datasets is detailed in Table 3. Owing to the effectiveness of multi-modality alignment, Babel consistently outperforms the SOTA methods for each individual modality. Notably, Babel demonstrates significant improvements for the mmWave and LiDAR modalities, achieving gains of 8.93% and 6.14%, respectively. In the Wi-Fi modality, Babel outperforms SenseFi by 4.4%. For the IMU modality, Babel attains an accuracy of 81.4%. It is important to note that none of the datasets evaluated here were included in the pre-training dataset collection, highlighting the generality of Babel.
Besides performance improvements, individual modalities could be empowered with new capabilities through Babel. For example, Wi-Fi sensing can be considerably improved in terms of its cross-domain capability. To showcase, we select two settings with different transmitter-receiver arrangements, denoted as S1 and S2 from the OPERANet dataset (Bocus et al., 2021). The encoder ViT (Dosovitskiy et al., 2020) trained on UT_HAR (Yousefi et al., 2017) attains an accuracy of 26% for classifying 6 activities on S2. This stands for the baseline of the cross-domain Wi-Fi sensing. Then we fine-tune the encoder on S1, then test on S2, achieving an accuracy of $54.5$%. It showcases the performance of the conventional domain adaptation approach. Finally, we use Babel, which accomplishes $62.47$% accuracy without any fine-tuning, evidencing its cross-domain capability.
Table 3. HAR classification accuracy of sensing modalities with Babel on four singular-modal datasets.
Modality HAR Classes Dataset Singular-modal Model Ours IMU 6 UCI (Reyes-Ortiz et al., 2016) LIMU-BERT (Xu et al., 2021) @ $80$% 81.4% Wi-Fi 22 Widar3.0 (Zhang et al., 2022) SenseFi (Yang et al., 2023a) @ $27.61$% 32.01% mmWave 12 mRI (An et al., 2022a) MARS (An and Ogras, 2021) @ $67.86$% 76.79% LiDAR 20 MSRAction3D (Li et al., 2010) Point Transformer (Zhao et al., 2020) @ $50.88$% 57.02%
@@
Cosmo is the SOTA sensing fusion framework, but unlike Babel, it requires all modalities to coexist within one dataset, limiting its expandability to datasets with complete paired data. Thus to compared with Cosmo, we utilize the same paired IMU-skeleton from the UTD-MHAD (Chen et al., 2015) that it excels. An equal amount of data is employed to train both Cosmo and a bi-modality version of Babel. We train Cosmo and Babel for 10 times with different random seeds. In the fusion of IMU and skeleton modalities, Cosmo achieves an averaged classification accuracy of 56.3%, while Babel attains 63.02% on UTD-MHAD.
What’s more, Cosmo’s performance can also be attributed to the integration of an additional network structure and its corresponding training procedure (for each downstream task), referred to as iterative fusion learning. Conversely, we aim to highlight Babel’s efficacy as a pre-trained network with a simple downstream task design. When applying the same downstream network (i.e., MLP), Babel could achieve around 22% accuracy improvement compared to Cosmo. Furthermore, as a expandable solution, Babel allows aligning more modalities without retraining pre-existing ones. This enhances Babel’s performance when introducing modalities.
@@
MLLMs IMU Video Wi-Fi mmWave Skeleton LiDAR UTD-MHAD (Chen et al., 2015) UTD-MHAD OPERANet (Bocus et al., 2021) XRF55 (Wang et al., 2024) UTD-MHAD MM-Fi (Yang et al., 2023c) OneLLM (Han et al., 2023; Zhang et al., 2023a) $6.5$% $6.51$% -- -- -- -- M4 (Yuan et al., 2024; Girdhar et al., 2023) $5.77$% $7.44$% -- -- -- -- Ours 31.77% 21.35% 33.89% 50.30% 61.06% 43.91%
@@
There has been a significant development in MLLMs (Han et al., 2023; Zhang et al., 2023a). These models are capable of understanding multi-modal inputs, including sensing modalities like IMU potentially. For comparison, we select typical MLLMs e.g., OneLLM (Han et al., 2023)and M4 (Yuan et al., 2024), and evaluate their performance on the UTD-MHAD (Chen et al., 2015) with HAR tasks. OneLLM and M4 use Meta-Transformer (Zhang et al., 2023a) and ImageBind (Girdhar et al., 2023) to interpret sensing signals, respectively. The results are summarized in Table 4. Firstly, current MLLMs can only support a limited number of sensing modalities, like IMU. Secondly, they only achieve a classification accuracy of around 5%-6%. In stark contrast, Babel significantly outperforms these with a classification accuracy of 31.77% on IMU while supporting other five sensing modalities.

### v2

@@ 3. Babel Overview
The crux of this design lies in the efficient alignment of representations across pre-trained encoders. Thereby, we introduce the modality tower, consisting of the pre-trained encoder, and the concept alignment module. The encoder could be based on signal processing and neural networks from existing deep learning models. The concept alignment module then aligns embeddings (features) from encoders. During training, pre-trained encoders are frozen, and the concept alignment module is updated.
In the design of the expandable model architecture, we try to convert the contrastive training process with N-tuple samples into a sequence of training phases involving only paired samples, thereby reducing the need for tupled samples, rendering the alignment of multiple modalities truly feasible.
@@ 4.1. Assembling Modality Towers
Comparing with conventional modality alignment methods i.e., CLIP, Babel’s key design lies in the utilization of a pre-trained encoder within a singular modality, proving particularly effective for sensing modalities. Sensing modalities (e.g., IMU, LiDAR, Wi-Fi) have matured over decades of research, leading to specialized feature extractors. These encoders incorporate domain knowledge in their architecture, training scheme, or signal-processing pipeline. By reusing these encoders rather than designing new ones from scratch, Babel benefits from the high-quality, domain-specific representations already learned in prior work.
@@
The particular encoder for each modality is chosen based on the following criteria. For modalities dedicated to sensing tasks, such as mmWave, we tend to choose signal processing-based encoders, due to their capability to extract universally applicable features with well-defined physical meanings. For more ubiquitous sensing modalities, like Wi-Fi, which are often noisy, we lean towards deep learning (DL)-based encoders, owing to their proficiency in de-noising. We avoid choosing models whose pre-training corpus might exhibit excessive domain shift compared to typical multi-modal sensing tasks. For example, a LiDAR encoder trained solely on autonomous-driving road scenes might not generalize well to indoor human-tracking data. Conversely, encoders trained on more general sets (e.g., broad 3D shapes for LiDAR, a diverse set of action videos) yield a feature space that better suits the variety of tasks Babel must handle. For modalities with large variation or especially noisy signals (e.g., Wi-Fi), relying on a single pre-trained encoder can be limiting. In such cases, we introduce modality tower augmentation, where multiple encoders are employed for the same modality, as elaborated in §4.3. Eventually we evaluate and compare the selected candidates by fine-tuning and testing them on a variety of singular modality datasets. The encoder demonstrating superior generality is chosen. To align modality embeddings, we employ a MLP layer as the concept alignment module. MLP-based projection layers have been widely used in contrastive learning frameworks (Chen et al., 2020).
Figure 3. The alignment of two modality towers with pre-trained encoders and concept alignment modules, where $e^{m}_{\alpha}$ denotes the $m$-th embedding of modality $\alpha$.
@@ 4.2. Aligning Modality Towers
where $L_{\alpha\leftarrow\beta}^{M}$ and $L_{\beta\leftarrow\alpha}^{M}$ denote the computed contrastive loss transitioning from modality $\beta$ to modality $\alpha$ and vice versa within the batch $M$, as defined subsequently,
(4) $L_{\alpha\leftarrow\beta}^{M}=-\sum_{i=1}^{m}\log\left(\frac{\exp(\text{sim}(P_{\alpha}^{i})/\tau)}{\sum\exp(\text{sim}(N_{\alpha}^{i})/\tau)}\right),$
@@ 5.2. Growth Orders
Babel transforms the $N$-tuple modality alignment into a sequence of two-modality alignment phases, thereby raising a potential question regarding the differences between the conventional completed alignment and our expandable alignment with varying modality growth orders. The insight of our prototype network is that it maintains a shared set of parameters across all modalities, inherently encoding common features learned from previous alignment phases. When aligning a new branch modality with the junction, the prototype network is updated in a way that partially shifts the shared embedding space, but does not overwrite it completely. This mitigates catastrophic forgetting (French, 1999), a prominent challenge in continual learning, which is usually addressed by utilizing shared representations to preserve previously learned information (Pham et al., 2021; Mallya and Lazebnik, 2018).
To analyze this, we take a three-modality alignment, i.e., IMU, skeleton, and video from UTD-MHAD dataset (Chen et al., 2015), as an example. As depicted in Fig.5, we utilize t-SNE to render the representation space of each modality visible. As evident in Figure 5(a), before alignment, features that have not undergone alignment training exhibit significant distribution differences. Fig. 5(b) shows the conventional triplet alignment successfully bridge the modality gaps, aligning the three modalities. In contrast, the expandable network architecture within Babel employs a sequence of two-modality alignment training phases as a replacement for the joint alignment. As illustrated in Fig. 5(c), we initially align the IMU and skeleton modalities followed by the video modality, effectively bridging the modality gaps as well.
@@ 6. Adaptive Training Strategy
We further propose our training strategies to optimally integrate the insights derived from the newly aligned modality during network growth. Specifically, we implement two strategies for the training of the concept alignment module and the prototype network, respectively.
@@
Gradient magnitudes effectively indicate how each modality contributes to the alignment process. During network growth, small gradients in the junction modality tower prompts a higher weight, bringing the branch network nearer to the trunk. When the branch modality tower’s gradients are significant, the assigned weight speeds up the absorption of insights from the trunk network, ensuring alignment in the unified representation space. Our approach employs bidirectional contrastive learning for each pairwise alignment phase (e.g.,$\alpha\leftarrow\beta$ and $\beta\leftarrow\alpha$), but rather than maintaining fixed equal weights between these directions—which can lead to either insufficient adaptation of the branch modality or excessive perturbation of the trunk network—we dynamically adjust the update magnitudes for both modules by monitoring their gradient norms during training. When the branch exhibits larger gradients, indicating that the branch modality is farther from alignment, we increase the weight on the junction modality (pulling the new modality “in”). Conversely, if the junction modality’s gradient is large, this suggests that the new modality is relatively close, allowing the junction side to shift more than at the beginning. This adaptive weighting scheme naturally preserves previously learned representations and, through sufficient training iterations, ensures convergence to a consistent common representation, regardless of the order in which modalities are incorporated.
@@ 7.1. Data Preparation
Overall, we utilize five datasets’ training sets for the alignment, as itemized in Table 1, comprising paired samples across divergent dual modalities. These datasets are for human activity recognition (HAR) tasks, but the certain activities are totally different. Despite the provision of activity labels within these datasets, we adopt a self-supervised training approach, labels are not used. Throughout each dataset, we split into train and test set, detailed in Table 1.
@@
Dataset Modalities # Train Pairs # Test Pairs
UTD-MHAD (Chen et al., 2015) IMU and Skeleton $613$ $248$
MM-Fi (Yang et al., 2023c) LiDAR and Video $17,528$ $3,132$
OPERANet (Bocus et al., 2021) Wi-Fi and Skeleton $25,433$ $5,086$
XRF55 (Wang et al., 2024) mmWave and Wi-Fi $30,000$ $12,900$
Kinetics-400 (Kay et al., 2017) Video and Skeleton $234,619$ $19,761$
Skeleton [^2] Depth signals undergo a conversion into a human skeleton format. As such, we employ the term skeleton to denote the depth modality and IMU pairs. UTD-MHAD dataset (Chen et al., 2015) is used, encompassing the skeleton and 9-axis IMU data pairs, captured via the Microsoft Kinect sensor and the wearable inertial sensor with respective sampling rates of 30Hz and 50Hz. The dataset embodies 27 distinct actions performed by 8 subjects. Each subject repeated the action for 4 times, totaling 861 paired samples. We use 613 pairs for the training.
@@
mmWave and Wi-Fi pairs. XRF55 dataset (Wang et al., 2024) is used, which is collected from a TI IWR6843ISK radar for mmWave and Intel 5300 for Wi-Fi CSI. It includes HAR data from 39 subjects performing 55 unique actions, each repeated 20 times. In total, 30,000 pairs are provided for training.
Video and skeleton pairs. Kinetics-400 dataset (Kay et al., 2017) is used, which contains 400 distinct human action classes, each characterized by at least 400 video clips extracted from YouTube. Each clip, approximately 10 seconds long, portrays a variety of human actions. The skeleton is extracted from use clips using OpenPose (Cao et al., 2017). Overall, as a dataset in vision modality, the Kinetics dataset provides 234,619 training pairs.
@@ 7.2. Data Augmentation
We implement two data augmentation techniques on the raw UTD-MHAD training data, ultimately enlarging the data pairs by 600$\times$. (i) Down-sampling. Raw pairs undergo down-sampling at different ratios, simulating diverse sampling rates on various devices or accelerating the action at distinct ratios. This method augments the raw pairs by a factor of 300$\times$. (ii) Action-segmentation. The raw action sequence is randomly truncated, simulating incomplete activity sensing. We ensure the segmented sequence’s shortest length is over 50% of the original length. This method amplifies the raw pairs by a factor of 300$\times$.
7.3. Pre-trained Encoders and Concept Alignment Architecture
Next we introduce the pre-trained encoders we use for building the modality alignment network, along with the concept alignment architecture.
@@ 7.3. Pre-trained Encoders and Concept Alignment Architecture
The concept alignment architecture comprises two Multi-Layer Perceptrons (MLPs) for each modality. This concept alignment module is customized for each modality, with the input dimensionalities determined by their encoders: LiDAR (60), Skeleton (60), IMU (72), Video (512), Wi-Fi (900), and mmWave (1024).
@@ 7.4. Training Details
We employ the AdamW optimizer (Loshchilov and Hutter, 2019) with a batch size of 256 and an initial learning rate of $1\times 10^{-4}$. Given a batch size of $m=256$, we construct 256 positive pairs and $65,280$ negative pairs for contrastive learning, following the construction method detailed in Section 4.2. For each phase of network growth, we judiciously allocate a varying number of training epochs, typically up to 500, or cease the training process once convergence is attained. The learning rate for downstream tasks is adjusted between 0.001 and 0.1. We train on two NVIDIA A100 GPUs, spending around 20 hours to align six modalities.
@@ 8. Evaluation
We evaluate pre-trained Babel by employing a typical downstream sensing task, human activity recognition (HAR). Furthermore, we would demonstrate two applications enabled by Babel: cross-modality retrieval and LLM integration.
@@ 8.1. Evaluation on HAR
We evaluate Babel on 8 datasets, comprising 4 in-domain datasets including UTD-MHAD (Chen et al., 2015), OPERANet (Bocus et al., 2021), XRF55 (Wang et al., 2024) and MM-Fi (Yang et al., 2023c). We utilize the test pairs outlined in Table 1. Additionally, we assess on 4 out-of-domain datasets, which were not part of the pre-training datasets at all: UCI (Reyes-Ortiz et al., 2016), Widar3.0 (Zhang et al., 2022), mRI (An et al., 2022a) and MSRAction3D (Li et al., 2010). Out-of-domain evaluation is critical for understanding model generalization in real-world sensing applications. In practical deployments, sensing models frequently encounter new environments and previously unseen subjects or activities. By evaluating on out-of-domain datasets, we can measure how well Babel adapts to new scenarios without any additional training.
The results for Babel are obtained in a one-shot setting, which serves as a widely used benchmark for evaluating the generalization capability of a backbone model (Fei-Fei et al., 2006; Koch et al., 2015). Unlike the conventional fully supervised scenario, one-shot learning is particularly relevant for sensing applications, where labeled samples are expensive and difficult to acquire (Xiao et al., 2021; Feng and Duarte, 2019). In this setup, we select only one labeled sample per class from the test portion for fine-tuning on the downstream task, while the remaining test samples are used to compute the accuracy metrics reported in our results tables. This challenging setting underscores Babel’s effectiveness as a pre-trained network, demonstrating its ability to generalize with minimal supervision in resource-constrained sensing environments.
@@
We compare Babel with a broad range of baselines, including SOTA singular-modal sensing baselines, LIMU-BERT (Xu et al., 2021) for IMU, SenseFi (Yang et al., 2023b) for Wi-Fi, MARS (An and Ogras, 2021) for mmWave, MeteorNet (Liu et al., 2019) together with PointTransformer (Zhao et al., 2020) for LiDAR. Additionally, we include the multi-modal sensing baseline Cosmo (Ouyang et al., 2022). Finally, we compare Babel with emerging MLLMs that hold potential for interpreting sensing signals, including OneLLM (Han et al., 2023) and M4 (Yuan et al., 2024).
(a) On in-domain datasets
(b) On out-of-domain datasets
Figure 7. One-shot HAR classification accuracy of each modality achieved by Babel on the in-domain and out-of-domain datasets compared to various singular-modal baselines: LIMU-BERT(IMU), SenseFi(Wi-Fi) (Yang et al., 2023a), MARS(mmWave) (An and Ogras, 2021), PointTransformer(LiDAR), ResNet3D(Video), ST-GCN(Skeleton). For out-of-domain datasets, Babel achieves 4.5×, 7.6×, 5×, and 10.4× improvements over random guessing (IMU: 16.7%, Wi-Fi: 4.5%, mmWave: 8.3%, LiDAR: 5.0%).
8.1.1. Performance on singular-modal sensing.
@@
Owing to the pre-trained alignment across multiple sensing modalities, Babel exhibits superior performance even when each individual aligned modality is applied to downstream tasks. In our evaluation, we apply one‐shot training and testing exclusively to each modality’s data—even if the dataset itself contains multiple modalities. For instance, although UTD‐MHAD includes IMU, video, and skeleton data for the same set of actions, we use only the IMU samples for one‐shot training and testing on the IMU, only the video samples for the video evaluation, and so forth. Fig. 7(a) presents the evaluation results of singular-modal sensing across four in-domain datasets. As illustrated, compared to SOTA singular-modal methods, Babel delivers an average accuracy improvement of approximately 12% across six aligned modalities in various datasets. Notably Babel achieves significant gains in weaker sensing modalities. For instance, classification accuracy for 27 human activities in the IMU modality increases from 20.19% to 31.77%, compared to LIMU-BERT on UTD-MHAD. The Wi-Fi modality obtains an approximate 10.74% enhancement. The mmWave modality shows a substantial increase from 30.32% to 50.30%, and the LiDAR modality achieves an accuracy of 43.91%, up from 28.43%. Such gains are achieved by aligning each modality into a unified representation space, facilitating mutual learning from strong modalities, i.e., video. Though sensing modalities benefit significantly, gains are limited for the video modality, increasing by around 2%.
@@
The performance of Babel on four full out-of-domain datasets is detailed in Fig. 7(b). Owing to the effectiveness of multi-modality alignment, Babel consistently outperforms the SOTA methods for each individual modality. Notably, Babel demonstrates significant improvements for the mmWave and LiDAR modalities, achieving gains of 14.6% and 13.1%, respectively. In the Wi-Fi modality, Babel outperforms SenseFi by 5.3%. For the IMU modality, Babel attains an accuracy of 74.9%. It is important to note that none of the datasets evaluated here were included in the pre-training dataset, highlighting the generality of Babel.
Table 2. The classification accuracy of sensing fusion for Babel and the multi-modal sensing baselines evaluated on the out-of-domain mRI(An et al., 2022a) dataset.
Multi-modal Baselines Babel Vision+IMU $89.6\%$ $91.7\%$(+$2.3\%$) Vision+mmWave $58.4\%$ $64.6\%$(+$10.7\%$) IMU+mmWave $75.0\%$ $86.5\%$(+$13.5\%$) Vision+IMU+mmWave $85.4\%$ $92.8\%$(+$8.6\%$)
8.1.2. Performance on multi-modal sensing.
The unified representation space in Babel allows for effective fusion. When IMU and Video modalities are fused, Babel achieves a 33.17% accuracy on UTD-MHAD (Chen et al., 2015), outperforming both the individual IMU and video modalities. Likewise, a 58.97% accuracy is achieved on XRF55 (Wang et al., 2024) when Wi-Fi and mmWave modalities are merged. Note that even modality combinations (like IMU&Video fusion) not included in pre-training datasets are evaluated and obtain the superior performance, highlighting Babel’s flexibility and offering developers numerous opportunities to choose any one or combined modalities for their tasks.
We also evaluate the multi-modal sensing fusion of Babel on an out-of-domain dataset, mRI(An et al., 2022a), which is not included in the pre-training datasets. Various combinations of modalities are evaluated, and Babel’s performance is compared with multi-modal sensing baselines. These baselines are implemented following the designs of existing works (An and Ogras, 2021; An et al., 2022b). As shown in Table 2, Babel achieves up to a 13.5% improvement in accuracy compared to the modality-specific baselines.
Table 3. Comparison with Cosmo (Ouyang et al., 2022) under various configurations for IMU-Skeleton fusion sensing: Babel (B) represents the Babel network that only aligns two modalities, while Cosmo(G) denotes the network customized its specific structure with simple MLPs.
Babel(B) Cosmo Cosmo(G) Babel Aligned Modalities 2 2 2 6 Downstream MLP Specific network MLP MLP Task Designs and training strategy IMU-Skeleton 61.46% 56.3% 41% 63.02% Fusion Acc.
@@
Cosmo is the SOTA sensing fusion framework, but unlike Babel, it requires all modalities to coexist within one dataset, limiting its expandability to datasets with complete paired data. Thus to compared with Cosmo, we utilize the same paired IMU-skeleton from the UTD-MHAD (Chen et al., 2015) that it excels. An equal amount of data is employed to train both Cosmo and a bi-modality version of Babel. We train Cosmo and Babel for 10 times with different random seeds. In the fusion of IMU and skeleton modalities, Cosmo achieves an averaged classification accuracy of 56.3%, while Babel attains 61.46% on UTD-MHAD, as shown in Table 3.
What’s more, Cosmo’s performance can also be attributed to the integration of an additional network structure and its corresponding training procedure (for each downstream task), referred to as iterative fusion learning. Conversely, we aim to highlight Babel’s efficacy as a pre-trained network with a simple downstream task design. When applying the same downstream network (i.e., MLP), Babel could achieve around 20% accuracy improvement compared to Cosmo. Furthermore, as a expandable solution, Babel allows aligning more modalities without retraining pre-existing ones. This enhances Babel’s performance when introducing modalities. As shown, Babel aligning six modalities during the pre-training phase could achieve 63.02% on the IMU-Skeleton fusion sensing task.
@@
MLLMs IMU Video Wi-Fi mmWave Skeleton LiDAR UTD-MHAD (Chen et al., 2015) UTD-MHAD OPERANet (Bocus et al., 2021) XRF55 (Wang et al., 2024) UTD-MHAD MM-Fi (Yang et al., 2023c) OneLLM (Han et al., 2023) with $6.5$% $6.51$% -- -- -- -- Meta-Transformer (Zhang et al., 2023a) M4 (Yuan et al., 2024) with $5.77$% $7.44$% -- -- -- -- ImageBind (Girdhar et al., 2023) Ours 31.77% 21.35% 33.89% 50.30% 61.06% 43.91%
@@
There has been a significant development in MLLMs (Han et al., 2023; Zhang et al., 2023a). These models are capable of understanding multi-modal inputs, including sensing modalities like IMU potentially. For comparison, we select typical MLLMs e.g., OneLLM (Han et al., 2023)and M4 (Yuan et al., 2024), and evaluate their performance on the UTD-MHAD (Chen et al., 2015) with HAR tasks. OneLLM uses Meta-Transformer (Zhang et al., 2023a) and M4 uses ImageBind (Girdhar et al., 2023) to interpret sensing signals, respectively. The results are summarized in Table 4. Firstly, current MLLMs can only support a limited number of sensing modalities, i.e., IMU. Secondly, they only achieve a classification accuracy of around 5%-6% according to our evaluation. In stark contrast, Babel significantly outperforms these with a classification accuracy of 31.77% on IMU while supporting other five sensing modalities.